### v1

# Object-level Geometric Structure Preserving for Natural Image Stitching

Abstract
The topic of stitching images with globally natural structures holds paramount significance. Current methodologies exhibit the ability to preserve local geometric structures, yet fall short in maintaining relationships between these geometric structures. In this paper, we endeavor to safeguard the overall, OBJect-level structures within images based on Global Similarity Prior, while concurrently mitigating distortion and ghosting artifacts with OBJ-GSP. Our approach leverages the Segment Anything Model [25] to extract geometric structures with semantic information, enhancing the algorithm’s ability to preserve objects in a manner that aligns more intuitively with human perception. We seek to identify spatial constraints that govern the relationships between various geometric boundaries. Recognizing that multiple geometric boundaries collectively define complete objects, we employ triangular meshes to safeguard not only individual geometric structures but also the overall shapes of objects within the images. Empirical evaluations across multiple image stitching datasets demonstrate that our method establishes a new state-of-the-art benchmark in image stitching. Our implementation and dataset is publicly available at https://github.com/RussRobin/OBJ-GSP
1 Introduction
Figure 1: Our panorama (d) preserves structures well even if the structure marked with red line is not recognized by SAM (b). By preserving the object-level structure of the building, the local shapes are also preserved.
Image stitching aims to align multiple images and create a composite image with a larger field of view. This method is widely utilized across diverse domains, including smartphone panoramic photography [43], robotic navigation [10], and virtual reality experiences [1, 24]. The field of image stitching algorithms grapples with two fundamental challenges: alignment and distortion. Autostitch [3] initially undertakes the extraction of feature points and computes the homography matrix to align feature points between images. Approaches such as APAP [47] and GSP [7] often segment the target image into smaller grids, computing local homographic transformations between the target and reference images within each grid. This, combined with global homography transformations, effectively addresses the alignment concern. However, a relentless pursuit of feature point alignment can inadvertently introduce distortion into the resulting images. In light of this, GES-GSP [13] commences with edge detection, extracting linear and curvilinear structures within the image. Subsequently, triangle sampling is applied to the curves. During the grid deformation process, consideration is given not only to the alignment of feature points within the grid but also to the preservation of similarity transformations between sampled triangles. This approach, thus, succeeds in safeguarding the geometric structures within the images. However, GES-GSP can only protect individual structures and lacks the ability to comprehend the relationships between structures at a global level. As shown in Fig. 1 (c), it fails to preserve the overall geometric structure.
In order to globally preserve the structure of objects and prevent distortion, we propose to leverage Segment Anything Model (SAM) [25] to extract object-level geometric information, and generate triangle meshes within each object to ensure the overall preservation of object shapes during the stitching process. SAM is capable of automatically segmenting common scenes and extracting semantically rich, high-quality, holistic geometric structures, as shown in Fig. 1 (b). SAM treats objects as cohesive entities, transcending the segmentation of individual line segments. Furthermore, SAM is proficient at extracting object outlines, surpassing the capabilities of conventional straight or curved lines and allowing for a more nuanced understanding of the relationships between geometric structures.
Our approach capitalizes on object-level geometric information. Drawing inspiration from rigid object deformations [19] in computer vision, we generate triangle meshes within each object. During the stitching process, these triangle meshes are collectively subjected to similarity transformations, effectively preserving the structure of the objects. As demonstrated in Fig. 1 (c) and (d), our method excels in maintaining the overall structure of objects, preventing distortion of prominent linear elements within the images, resulting in a more natural and faithful stitching outcome. Notably, even in scenarios where SAM fails to extract structures below the red dashed line, our approach remains effective by preserving the holistic structure of the object with triangle meshes.
Our proposed approach has resulted in a new state-of-the-art in the field of image stitching. The main contributions of this paper are as follows:
  • We introduce semantic information into GES-GSP [13] through SAM [25], enabling the extraction of semantically enriched, uniformly distributed object-level image structures.
  • We generate triangular meshes within objects and constrain them as a whole to undergo similarity transformations, preserving object shapes globally.
  • Through experiments on standard image stitching datasets, our method surpasses GES-GSP and establishes a new state-of-the-art benchmark.
2 Related works
Grid-based image stitching algorithm. Autostitch [3], a pioneering work in image stitching, identifies matching feature points and aligns them by homography transformation. Building upon this foundation, numerous stitching algorithms partition images into grids, compute geometric transformation relationships for each grid, and combine them into a global transformation to align overlapping regions and seamlessly transit the transformation to non-overlapping areas.
APAP [47] warps aim to maintain global projectivity while allowing local non-projective deviations to account for deviations from assumed imaging conditions. AANAP [28] linearizes the homography and gradually shifts it towards global similarity to mitigate perspective distortion. GSP [7] achieves more natural stitching results by leveraging local or global similarity transformations. However, grid deformation introduces some geometric distortion. To address this, LPC [20] proposed to leverage line-point consistency to preserve structures, introducing global collinear structures to enhance desirable characteristics for image warping. GES-GSP [13] not only preserves straight-line shapes but also protects the image’s curved structures, allowing curves to undergo similarity transformations to prevent distortion.
Nevertheless, LPC [20] and GES-GSP’s [13] extraction of straight or curved lines fails to fully capture the image’s geometric structure. Their structure preservation algorithms only protect individual lines without recognizing the overall structure. To address this, we employ semantic segmentation to extract meaningful geometric structures, use triangular grids to protect the overall object-level geometric structure, and establish connections between dispersed geometric transformations, achieving superior results.
Geometric structure extraction. In LPC [20] and GES-GSP [13], the Line Segment Detector (LSD) [40] algorithm detects straight lines in images. GES-GSP employs edge detection methods like Canny [4] and HED [42] to identify edges, which represent the geometric structure of curves. LSD efficiently detects line segments based on gradient information. Canny is an image processing technique that identifies edges by detecting significant changes in pixel intensity, known for its noise suppression and precise edge mapping capabilities. HED is a neural network-based end-to-end edge detection method that outperforms Canny in edge detection. In recent years, deep learning-based semantic segmentation [38, 37, 6, 18, 46, 8] and instance segmentation [16, 17, 2, 23] methods have advanced rapidly, capable of segmenting objects with semantic information in images. However, they require large datasets [9, 31, 39] for training and can only segment fixed categories within the dataset. Thus, large-scale pretrained encoders [5, 35, 36] have been proposed under the encoder-decoder architecture, capable of fine-tuning for downstream tasks. SAM [25], introduced in early 2023, is a large model capable of semantic segmentation for point, crop box, and text prompts. SAM’s automatic mode automatically selects point prompts, segments images, and extracts object information.
We employ SAM [25] to extract object-level geometric structures and preserve these shapes during stitching. Compared to LSD [40], Canny [4], and HED [42], SAM’s extracted geometric structures carry semantic meaning, distribute evenly, and have a pleasing visual appearance. LPC [20] and GES-GSP [13] only protect independent, scattered geometric lines, while we protect objects composed of multiple geometric lines as a whole, resulting in superior performance.
3 GES-GSP pipeline
GES-GSP [13] uses a grid mesh to guide the image deformation, where $V$ and $E$ represent the sets of vertices and edges within the grid mesh. GES-GSP aims to find a set of deformed vertex positions, denoted as $\widetilde{V}$, that minimizes the energy function $\psi(V)$. In summary, GES-GSP [13] builds upon the foundation of GSP [7] by introducing the GES term, specifically aimed at preserving the shape of lines, in addition to the Alignment term, local similarity term, and global similarity term within GSP [7]. In this section, we provide a brief overview of the optimization equations employed in GES-GSP [13].
Alignment term extracts feature points from the input images with the goal of aligning these points after geometric transformation. For each feature point pair $(p,\Phi(p))$, $\widetilde{v}(p)$ represents the position of $p$ as a linear combination of four vertex positions, and $M$ represents the set of all feature point pairs. The algorithm linearly combines the coordinates of the four vertices of each grid to represent the position of $p$ through bilinear interpolation. By optimizing the positions of grid vertices after geometric transformation, it aims to bring $p$ as close as possible to $\Phi(p)$. Therefore, the energy equation is defined as:
$\psi_{a}(V)=\sum_{p_{k}\in M}\|\widetilde{v}(p_{k})-\widetilde{v}(\Phi(p_{k}))\|^{2}.$ (1)
Local similarity term aims to ensure that the transition from overlapping to non-overlapping regions is natural and undistorted. Each grid undergoes a similarity transformation to minimize shape distortion. For an edge $(j,k)$, $S_{jk}$ represents its similarity transformation. Suppose $v_{j}$ transforms to $\widetilde{v_{j}}$ after deformation, and the energy function is defined as:
$\psi_{l}(V)=\sum_{(j,k)\in E_{i}}\|(\widetilde{v}_{k}-\widetilde{v}_{j})-S_{jk}(v_{k}-v_{j})\|^{2}.$ (2)
Global similarity term operates on a global scale to ensure the entire image undergoes a similarity transformation. GSP algorithm evaluates the scale $s$ and rotation $\theta$ within the global image transformation and computes parameters $c(e)$ and $s(e)$ for similarity. Thus, the energy function is defined as:
$\psi_{g}(V)=\sum_{e_{j}\in E}w(e_{j})^{2}[(c(e_{j})-s\cos\theta)^{2}+(s(e_{j})-s\sin\theta)^{2}].$ (3)
GES term in GES-GSP [13] involves extracting straight lines and curve structures from the image, followed by triangulation to obtain sampled points $\widehat{V_{i}}$. These points transform to ${\widehat{V_{i}}}^{desired}$ after projection transformation and similarity transformation. To bring $\widehat{V_{i}}$ as close as possible to ${\widehat{V_{i}}}^{desired}$ after the final grid deformation, GES-GSP defines the energy term as:
$\psi_{ges}(V)=\sum_{\beta=1}^{N_{c}}\sum_{\alpha=1}^{N_{s}}{\omega_{\alpha}^{\beta}E_{\alpha}^{\beta}}.$ (4)
For a detailed description of the four energy terms mentioned above, please refer to GSP and GES-GSP for more information. Ultimately, the sought-after vertices are determined by minimizing the following optimization equation for two images:
$\widetilde{V}=\text{argmin}_{\widetilde{V}}(\psi_{a}(\widetilde{V})+\lambda_{l}\psi_{l}(\widetilde{V})+\psi_{g}(\widetilde{V})+\lambda_{ges}\psi_{ges}(\widetilde{V})).$ (5)
The above equation represents the optimization for two images. For a scenario involving $n$ images, such as $I_{0}$, $I_{1}$, $I_{2}$, up to $I_{n-1}$, with $I_{0}$ as the reference image and the remaining $n-1$ images transformed onto the plane of $I_{0}$, the corresponding energy terms are denoted as $\psi^{1}$, $\psi^{2}$, and so on up to $\psi^{n-1}$. The final optimization is performed as:
$\widetilde{V}=\text{argmin}_{\widetilde{V}}(\psi^{1}+\psi^{2}+\ldots+\psi^{n-1}).$ (6)
4 Proposed method
In GES-GSP [13], it is necessary to extract large-scale geometric structures, including straight lines and curves, from the entire image and preserve these structures during stitching. Therefore, GES-GSP employs an edge detection algorithm to extract edges from the image. Subsequently, the image undergoes skeletonization, noise reduction, branch pruning, and corner point removal operations. Afterward, the LSD [40] algorithm is applied to detect straight lines within the image. Post-processing techniques are then employed to handle invalid edge information, including branch pruning, reconnection of discontinuous lines, and collinearity constraints. While this conventional approach can extract some geometric structures, its results often appear meaningless to human observers and fail to capture the complete geometric information of the image.
To address this issue, our OBJ-GSP variant aims to preserve structures imbued with semantic information, making the protected structures more semantically meaningful. Moreover, the algorithm used in GES-GSP [13] tends to overemphasize certain regions, leading to an excessive weighting of these areas during the stitching process. To mitigate this, we adopt semantic segmentation methods to extract object-level geometric structures. These structures are uniformly distributed and clearly delineate the boundaries of objects in the image, enabling the preservation of meaningful geometric information in the stitched result.
Figure 2: (a) Our triangle mesh, $V_{0}$ being the center of object. (b) With every edge undergoing similarity and projection transformation, object in (a) is transformed into (b). (c) Triangle mesh with near-equilateral triangles of similar sizes across the region. (d) Triangle sampling strategy employed in GES-GSP.
After obtaining the contours, we generate a triangular mesh for each semantic object, preserving the shape of the object through similarity transformations within the triangle mesh. Unlike the As-Rigid-As-Possible (ARAP) [19] method, we simplify computational complexity by directly locating the center of the object and connecting it to sampling points on the object’s semantic boundary to form a triangular mesh. In Fig 2, $V_{0}$ represents the object’s center, while $V_{1}$ and $V_{2}$ are sampling points on the semantic boundary of the object, forming a triangular mesh with these three points. One vertex, $V_{2}$, of the triangle can be represented using the edges of the triangle $\vec{V_{0}V_{1}}$ and an orthogonal coordinate system obtained by rotating this edge counterclockwise by 90 degrees:
$V_{2}=V_{0}+x_{01}\overrightarrow{V_{0}V_{1}}+y_{01}\left[\begin{matrix}0&1\\ -1&0\\ \end{matrix}\right]\overrightarrow{V_{0}V_{1}}.$
After the mesh deformation, $V_{0}$ and $V_{1}$ are transformed into $\widehat{V_{0}}$ and $\widehat{V_{1}}$. To preserve the shape of the segmentation result, we aim for the triangle to undergo a similarity transformation, keeping $x_{01}$ and $y_{01}$ unchanged. Therefore, we desire $V_{2}$ to transform into:
$\widehat{V_{2}^{desired}}=\widehat{V_{0}}+x_{01}\overrightarrow{\widehat{V_{0}}\widehat{V_{1}}}+y_{01}\left[\begin{matrix}0&1\\ -1&0\\ \end{matrix}\right]\overrightarrow{\widehat{V_{0}}\widehat{V_{1}}}.$
The corresponding energy term for the transformed $\widehat{V_{2}}$ is calculated as:
$E_{V_{2}}=\|\widehat{V_{2}^{desired}}-\widehat{V_{2}}\|^{2}.$
Similar definitions for energy terms are applied to $\widehat{V_{0}}$ and $\widehat{V_{1}}$, resulting in the error sum for a triangle:
$E_{\{V_{0},V_{1},V_{2}\}}=\sum_{i=0,1,2}\|\widehat{V_{i}^{desired}}-\widehat{V_{i}}\|^{2}.$
Initially, our approach constructs the triangular mesh by selecting sampling points and the object’s center. Unlike ARAP [19], we do not employ equilateral triangular meshes, as objects segmented from the image often lead to very small equilateral triangles. Experimental results demonstrate that this approximation not only has no adverse impact on the final outcome but also reduces computational complexity:
$E_{{{V_{0},V_{1},V}_{2}}}=\sum_{i=1,2}\|\widehat{V_{i}^{desired}}-\widehat{V_{i}}\|^{2}.$
We extract $N_{c}$ semantic object structures from a single image using semantic segmentation, and $N_{s}$ represents the total number of all sampling points within geometric structure i. Similar to GES-GSP [13], $\omega$ is a coefficient calculated based on the positions of the sampling points. Consequently, the total error equation is as follows:
$\psi_{ges-o}(V)=\sum_{\beta=1}^{N_{c}}\sum_{\alpha=1}^{N_{s}}{\omega_{\alpha}^{\beta}E_{\alpha}^{\beta}}.$ (7)
Ultimately, our objective function is given by:
$\begin{split}\widetilde{V}&=\underset{\widetilde{V}}{arg\min}\left(\psi_{a}(\widetilde{V})+\lambda_{l}\psi_{l}(\widetilde{V})\right.\\ &\quad\left.+\psi_{g}(\widetilde{V})+\lambda_{ges-o}\psi_{ges-o}(\widetilde{V})\right).\end{split}$ (8)
For fair comparison, our parameters are identical to those of GES-GSP: $\lambda_{l}=0.75$, $\lambda_{ges-o}=1.5$. Our $\lambda_{ges-o}$ corresponds $\lambda_{ges}$ in to GES-GSP.
Models / Datasets AANAP [28] APAP [47] CAVE [22] DFW [27] DHW [14] GES-GSP [13]
Mean Distorted Residuals (MDR)
GSP [7] ECCV’16 1.06183 1.25495 0.90884 0.98457 1.08755 1.06986
GES-GSP [13] CVPR’23 1.06213 1.24249 0.90821 0.98034 1.05619 1.04473
OBJ-GSP (ours) 1.05930 1.20123 0.89731 0.97259 1.00496 0.98288
Improvement(%) 0.3 3.3 1.2 0.8 4.9 5.9
Naturalness Image Quality Evaluator (NIQE)
GSP 2.84241 3.4356 4.04708 5.61905 2.75485 3.84897
GES-GSP 2.77220 3.48713 4.03835 5.71544 2.70838 3.79240
OBJ-GSP (ours) 2.74965 3.39280 4.01565 5.69104 2.60825 3.70041
Improvement(%) 0.8 2.7 0.6 0.4 3.7 2.4
Models / Datasets LPC [20] REW [26] SEAGULL [30] SVA [32] SPHP [29] Our dataset
Mean Distorted Residuals (MDR)
GSP 1.30562 1.16192 1.14467 1.51158 1.17784 1.15296
GES-GSP 1.22256 1.15462 1.11993 1.47197 1.14340 1.14366
OBJ-GSP (ours) 1.10622 1.08635 1.08296 1.47813 1.07699 1.12229
Improvement(%) 9.5 5.9 3.3 -0.4 5.8 1.9
Naturalness Image Quality Evaluator (NIQE)
GSP 4.28546 3.18549 5.10784 7.00495 3.04781 2.66597
GES-GSP 3.35315 2.75234 4.69390 6.96670 2.97173 2.64986
OBJ-GSP (ours) 3.23057 2.81480 4.08903 6.96149 2.49712 2.54906
Improvement(%) 3.7 -2.2 12.9 0.5 16.0 3.8
Table 1: Quantitative results on 9 image stitching datasets. We report the mean MDR and NIQE on each dataset. We label best results with bold text. A lower MDR and NIQE indicates better stitched panorama. Improvement row compares our proposed method and GES-GSP. The whole dataset consists of 124 images and our mean improvement to GES-GSP is 3.5% in MDR and 3.8% in NIQE.
5 Experiments
Figure 3: APAP and SPHP see misalighment, and we delineated the indistinct portions using color-coded boxes. Autostitch, APAP, SPHP and GSP sees distortion. The convergence of the blue and red lines is essential, and we signify distortion by the intersection of these two lines. GES-GSP successfully prevents distortion, but it undergoes misalignment. Our method addresses misalignment and distortion well.
Figure 4: Sampling points in our OBJ-GSP focus more on main structures so we can stitch precisely, as shown in (b)(d).
In this section, we conduct a quantitative comparison among GSP [7], GES-GSP [13], and our proposed OBJ-GSP. Additionally, we qualitatively compared our results with those obtained from other image stitching methods, such as Autostitch [3], APAP [47] and SPHP [29], through visual assessment of the stitched images.
5.1 Experimental setup
Evaluation metrics. We assess the quality of our stitching results from two perspectives. First, we employ the Mean Distorted Residuals (MDR) metric introduced in GES-GSP [13] to measure the degree of image distortion. In intuitive terms, if the points on the same side of the mesh were originally collinear and remain collinear after stitching, it implies that the stitching result has minimal distortion. Please refer to supplementary material of [13] for more details of MDR. Furthermore, we employ the Naturalness Image Quality Evaluator (NIQE) [33] metric to evaluate image quality and clarity. Compared to other image quality assessment methods like Structure Similarity Index Measure (SSIM) [41] and Peak Signal-to-Noise Ratio (PSNR), NIQE does not require a reference image. Since there is no strict ground truth for image stitching, we use NIQE to assess the quality of our stitching results. However, it is essential to emphasize that in the field of image stitching, there is no universally accepted objective evaluation method for assessing the quality of stitching results or the degree of distortion. Many practitioners still rely on subjective assessments based on human visual perception to gauge the quality of stitching outcomes.
Datasets. We utilize datasets commonly associated with previous image stitching algorithms, which include datasets sourced from AANAP [28], APAP [47], CAVE [22], DFW [27], DHW [14], GES-GSP [13], LPC [20], SEAGULL [30], REW [26], SVA[32] and SPHP [29]. We also collected our own dataset of 18 image pairs for image stitching. Thus the datasets used in experiments encompass a diverse range of settings, including indoor and outdoor environments, various regions around the world, urban and natural landscapes, among others. Please kindly refer to the supplementary material for details on how we collected our own dataset and conducted an analysis of the ten datasets we use.
Implementation details. We evaluated the results of GSP [7] and GES-GSP [13] using their publicly available C++ code. Our stitching code is also implemented in C++. Our implementation of SAM includes two approaches. Drawing from the Segment Anything C++ Wrapper [11], we exported the SAM [25] encoder and decoder into Open Neural Network Exchange (ONNX) format and subsequently replicated SAM’s automatic mode within C++. To achieve the best possible stitching results, we also directly implemented the semantic segmentation component using the SAM’s publicly available Python code and utilized their semantic segmentation results. The stitching part of our experiments ran on the CPU, while the SAM modules were capable of running on both CPU and GPU. Please kindly refer to the supplementary material on how we re-establish GSP and GES-GSP baselines.
5.2 Quantitative results
Table 1 shows MDR and NIQE results on datasets used in other stitching algorithms and our own dataset. The lower MDR and NIQE of OBJ-GSP indicates better stitching capability.
5.3 Qualitative results
In the context of image stitching, human perception is primarily influenced by two key factors: alignment and distortion. Misalignment can result in undesirable artifacts such as ghosting and blurriness, while distortion can make the stitched image appear unnatural. Fig. 3 llustrates the stitching outcomes of six different methods, where we use straight lines and boxes to demonstrate the effects of alignment and distortion. GES-GSP addresses the issue of distortion in the stitching results, and building upon this foundation, our OBJ-GSP method achieves superior alignment performance. Fig. 1 and 4 elucidates the reasons behind the superior performance of our OBJ-GSP method. With the assistance of semantic segmentation techniques, we place greater emphasis on preserving critical structures and ensuring holistic protection at the object-level for objects within the images. Please kindly refer to our supplementary material for more results.
6 Ablation Studies
Semantic segmentation methods. To assess the influence of semantic segmentation results on the stitching performance, we conducted a comparative analysis across three different backbones of the SAM [25] model, namely ViT-B, ViT-L, and ViT-H [12], with a progression from smaller to larger models. Larger models are inherently capable of capturing more fine-grained semantic details. The corresponding stitching results are depicted in Figure 5. It is worth noting that the models based on ViT-B and ViT-L exhibit some blurriness and minor distortions. From the perspective of MDR, our model, under the three aforementioned backbone configurations, achieved improvements of 4.1%, 8.3%, and 9.4% over the baseline model GES-GSP [13], respectively.
Figure 5: GES-GSP and the proposed OBJ-GSP with three semantic segmentation backbones.
Sampling strategies. Incorporating the SAM [25] with ViT-H [12] backbone, the utilization of triangular grid sampling yields superior results compared to the triangular sampling proposed in GES-GSP [13]. The triangular grid preserves the shape of lines but fails to maintain the overall geometric structure of the image. As illustrated in the Fig. 6, we have outlined the structural elements in the original image using red dashed lines and then superimposed them onto the combined results of triangular sampling and triangular grid sampling. Triangular grid sampling retains the positional relationships between the lines present in the original image.
Figure 6: From top to down: one of the input images, OBJ-GSP with triangle samping and with triangle mesh sampling.
7 Discussion
Limitations. While our approach has demonstrated state-of-the-art performance in image stitching by extracting object-level geometric structures with SAM and preserving them with triangle mesh sampling, it still has several limitations:
1. Our method involves the use of large models for semantic segmentation, which results in longer processing times. According to our analysis, the effectiveness of geometric structure extraction significantly impacts the final results. Our method is constrained by the quality of SAM’s results. Smaller models like SAM [25] with Vit-B/L [12] do not perform as well as SAM with Vit-H. Fig. 5 demonstrates the differences.
2. As shown in Fig. 7, the proposed method fails in cases of significant parallax. In the perspective of the left image, the corner of the building appears to be obtuse. However, based on the inference from the right image, the corner of the building should be a right angle. Consequently, the stitching algorithm becomes perplexed, unsure of how to preserve the shape of the house. As discussed in [3], in an ideal scenario without parallax, one would only rotate the camera for capturing images without changing the position of the camera’s optical center. In general, we have not observed a significant improvement compared to GES-GSP in Fig. 7.
In only rare instances does OBJ-GSP experience failure due to the malfunction of SAM [25], such as in conditions characterized by inadequate illumination or an exceedingly sparse set of features. Please refer to the supplementary material for details.
Figure 7: Failure case. Preserved structures are marked with red lines, while structures marked with blue lines are distorted due to considerable disparity in their perspectives within the input images. Although OBJ-GSP successfully preserve structures in green box, we see no apparent overall improvement over GES-GSP.
Seam-driven and deep learning based image stitching approaches. This paper primarily focuses on grid-based image stitching methods, which are currently the most explored and popular techniques. Previous work has explored seam-driven image stitching [15]. Seam-based methods, often combined with image blending [44, 45], have been widely used for post-processing in image stitching. The stitching approach discussed in this paper alleviates the need for extensive post-processing, resulting in more natural stitching results. As in provious grid-based works, we do not compare OBJ-GSP with seam-driven methods, but OBJ-GSP can be integrated with these post-processing methods. In recent years, several methods [21, 34] have attempted to model certain image stitching steps as unsupervised deep learning problems, leading to notable advances in the field.
Broader impacts. It’s important to acknowledge that our paper doesn’t explicitly discuss broader impacts in the proposed OBJ-GSP image stitching algorithm, such as fairness or bias. Segment Anything Model (SAM) [25] has discussed its broader impacts regarding geographic and income representation, as well as Fairness in segmenting people. Further research into how our algorithm may interact with other aspects of image stitching is encouraged.
8 Conclusion
In this paper, we propose Object-level geometric structure preserving for natural image stiching (OBJ-GSP) algorithm, which stands as a novel approach to achieving natural and visually pleasing composite images. The existing state-of-the-art, GES-GSP [13], while effective, faced limitations in accurately capturing geometric structures and preserving spatial relationships between these structures. In response, OBJ-GSP leverages semantic segmentation techniques, specifically the Segment Anything Model (SAM) [25], to extract object-level geometric structures imbued with semantic information. The proposed OBJ-GSP method addresses the shortcomings of its predecessor by preserving structures that align more intuitively with human perception. It employs triangular mesh deformations within individual objects to safeguard not only individual geometric structures but also the overall shapes of objects within the images. Empirical evaluations across multiple image stitching datasets have unequivocally demonstrated that OBJ-GSP surpasses the performance of GES-GSP, establishing a new state-of-the-art benchmark in image stitching. In conclusion, the OBJ-GSP algorithm offers a promising direction for image stitching, enhancing the preservation of geometric structures and object-level details, ultimately leading to more natural and visually appealing composite images.
[…]
Supplementary Material
We would like to clearly show some experiment and dataset details in this supplementary material. In the supplementary material, we provide more qualitative results, and in depth discussion and results on why Segment Anything Model [25] helps the proposed OBJ-GSP image stitching model, and failures cases of SAM.
9 Dataset
We employed handheld smartphones to capture images. During the image acquisition process, we took care to minimize translational movement of the smartphone, primarily relying on rotation to adjust the framing. This approach was employed to ensure that the disparity between images remained manageable, preventing situations where the occlusion relationships between two images would be too dissimilar for successful image stitching. We amassed a total of 18 pairs of image sets, encompassing diverse scenes such as rooms, culinary creations, sculptures, gardens, rivers, ponds, industrial facilities, roads, and exteriors of buildings, among others. Additionally, we curated existing image stitching datasets [28, 47, 22, 27, 14, 13, 20, 30, 26, 32, 29] to supplement our data collection efforts. Ultimately, we constructed a dataset consisting of 124 groups of images, marking it as the largest dataset currently employed in image stitching endeavors. We intend to release this dataset to the public for further research and development.
10 Reestablishing baselines
We replicated the results reported in the literature for GSP [7], GES-GSP [13], APAP [47], and SPHP [29] by implementing their publicly available codebases with their default parameter settings. For the structural alignment component, we employed the executable provided by Autostitch [3]. GES-GSP includes experimental data for both GES-GSP and GSP. In our experiments, the performance results for GES-GSP were consistent with the reported data, but GSP exhibited notably superior performance. This discrepancy may be attributed to variations in parameter configurations, as GES-GSP did not disclose the specific parameter settings used for the GSP experiments. In our method’s experiments, we maintained consistent parameter settings across all trials. Furthermore, for the structure extraction stage, we utilized the official code provided by SAM. However, we excluded masks with extremely small areas from our analysis.
11 More qualitative results
In this section we provide more qualivative results. We mark images with boxes to indicate misalignment, and use lines (and intersections of lines) to show distortion. Please refer to Fig. 8, 9, 10, 11 and 12. It is shown that our object-level preservation of structures can prevent distortion and misalignment at the same time.
12 Successful cases because of Segment Anything Model
In this section, we demonstrate the superiority of Segment Anything Model [25] with qualitative results. Please refer to Fig. 13 It is shown that SAM extracted object-level and complete structures of the ground and mountain, so the OBJ-GSP preserved their structures better than GES-GSP, where HED [42] only extracted fragmented edge information. In Fig. 14, we demonstrate that the proposed OBJ-GSP can even stitch images with poor lighting conditions and protect their object-level structures.
13 Failure cases because of Segment Anything Model
We acknowledge that it’s very important to discuss the situations where, Segment Anything Model [25] fails, leading to the failure of the proposed OBJ-GSP image stitching algorithm. Please refer to Fig. 15 and 16. In the context of our approach, it is not imperative for the SAM [25] to achieve precise segmentation of objects containing semantic information. SAM need only recognize key object contours and boundaries. Therefore, the proposed OBJ-GSP is susceptible to SAM failure only in exceedingly rare instances where features are exceptionally sparse, and objects are highly indistinct. Consequently, SAM’s failure would impact solely the maintenance term of our structure, leading OBJ-GSP to degrade to the performance level of the GSP [7]. Simultaneously, we emphasize that SAM’s failure would result in the nullification of our structural preservation term, causing OBJ-GSP’s performance to regress to that of GSP only under scenarios where SAM proves ineffective. In addressing cases involving distortion and misalignment, we posit that mitigation strategies such as global straightening and multi-bend blending, as employed in Autostitch [3], can be leveraged to alleviate these issues.
Figure 8: More qualitative results. Images are from LPC [20] dataset. As marked with the red line, Autostitch, SPHP and GES-GSP are distorted during transformation. APAP and GES-GSP sees misalignment. OBJ-GSP protects the shape of river bank and precisely aligns input images.
Figure 9: More qualitative results. Images are from LPC [20] dataset. Almost all methods except from GSP and OBJ-GSP misalign images. GSP distorts the shape of building during transformation. OBJ-GSP tries to reach a balance between shape preservation and alignment, although there’s some slight distortion in the green box.
Figure 10: More qualitative results. Images are from REW [26] dataset. We magnify the ground in the red box and the car in the purple box. OBJ-GSP aligns well but the remaining five methods shows misalignment. Distortion is not observed in this case.
Figure 11: More qualitative results. Images are from SPHP [29] dataset. GSP and GES-GSP suffers from misalignment as shown in the red boxes. The misalignment of Autostitch and APAP are shown in purple boxes. Also, SPHP distorts the bridge and ground.
Figure 12: More qualitative results. Images are from GES-GSP [13] dataset. Misalignment can be seen in results of Autostitch, APAP and SPHP. GES-GSP also distorts images. The stitched result of GSP is stretched. In contrast, the proposed OBJ-GSP aligns images well and protects pbject structures.
Figure 13: SAM successful case. Compared to HED [42], SAM extracted complete structures of the mountain and ground. As a result, OBJ-GSP can protect their structures better, although the left part of the ground is slightly distorted.
Figure 14: SAM successful case. Thanks to structures extracted by SAM, OBJ-GSP is able to stitch images taken at nights, with limited features. OBJ-GSP protects the structure of trees, as marked with red lines, because SAM’s structures are more informative than HED’s [42].
Figure 15: SAM failure case. The stitched pano of OBJ-GSP is distorted and misaligned because of the failure of SAM, but it can be metigated by global straightening and multi-bend blending.
Figure 16: SAM failure case. SAM fails to extract meaningful object-level structures, so OBJ-GSP performs at the same level of GSP [7].